OpenStudy (anonymous):

i will give medals Explain, in complete sentences, how you would use the substitution method to solve the following system of equations. Provide the solution to the system. 5x - y = -1 7x - 3y = 13

OpenStudy (anonymous):

medal really? 1) In the first equation, isolate y on one side of the equation and all other terms on the other. 2) substitute the value for Y determine in step 1 into equation 2 every place that Y appears. there is now an equation of a single variable, X. 3) solve for X, which gives a numeric value. 4) substitute that value for X back into either original equation to solve for Y. the final answer is x=-2, y=-9

OpenStudy (anonymous):

ok i'm ready for my awesome medal

OpenStudy (anonymous):

first put y interms of x. that is y=5x+1,then put it in equation two that is 7x-3(5x+1) = 13. then 7x-15x-3 =13 then -8x = 16 which is x = -2. now Y =5(-2) + 1 = -10 +1.which is Y = -9.